Coupons driving visits.

A store randomly samples 603 shoppers over the course of a year and finds that 142 of them made their visit because of a coupon they'd received in the mail. Construct a 95% confidence interval for the fraction of all shoppers during the year whose visit was because of a coupon they'd received in the mail.
Answer #1

Answer

Given that

sample size n = 603

number of shoppers made visit (x) = 142

sample proportion p(hat) = x/n = 142/603 = 0.2355

z critical = 1.96 (using z table)

confidence interval =